a car starting from rest has a speed of 30km/h at any one instant. two seconds later, its speed is 36km/hr and 2sec after that, it is 42km/hr, its accelaration in m/s^2 is?

Grade:11

1 Answers

Vikas TU
7 years ago
For fiest instant,
v = u + at
30 = 0 + a*t............(1)
After 2 seconds that is after (t+2)
36 = 30 + a(t+2)
or
6 = at + 2a..........(2)
solving both eqns we get,
2a = -24
a = -12 m/s^2.
which could not be possible as it has travelled further more after these 2 seconds but the accln is deccelation.
Hence not possible.